OpenStudy (deathslasher99):

Which rational number expression is larger? a. 2.75 b. 22/8 c. the numbers are equal

OpenStudy (calculusxy):

Convert 22/8 to decimal

OpenStudy (calculusxy):

\(22 \div 8 = 2.75\) Therefore, the are equal.

OpenStudy (calculusxy):

In questions like this, make sure to have both of the numbers in the same format. It will be easier for comparing which one is greater or lesser.

OpenStudy (kittiwitti1):

(just another method but calc's is easier) • • • You can also convert 2.75 to a mixed number and then improper fraction:\[2.75\rightarrow2\frac{3}{4}\]To convert, multiply the 2 by the 4, because \(2\rightarrow2\times\frac{1}{4}\), then add 3 because it's the remainder;\[\frac{(2\times4)+3}{4}\]This gets you the following:\[\frac{8+3}{4}=\frac{11}{4}\]Now, you need the fractions to have an equal denominator, so multiply the bottom by 2 to get 8. And since the top and bottom both need to be balanced:\[\frac{11}{4}\times\frac{2}{2}=\frac{11\times2}{4\times2}=\frac{22}{8}\]So yes, they are both equal.